OpenStudy (anonymous):

Find the unknown side as well as the perimeter of the triangle. express your result to the nearest tenth of a foot. left sided triangle shape measured at 2 ft, with the base at 6 ft.

OpenStudy (dumbcow):

is this a right triangle?

OpenStudy (anonymous):

(for a right triangle only) \[a^{2}+b^{2}=c^{2}\] \[2^{2}+6^{2}=c^{2}\] \[44^{2}=c^{2}\] \[\sqrt{44^{2}}=\sqrt{c^{2}}\] 6.63=c (or the unknown side) P (perimeter) = 2+6+6.63 P=14.63ft

OpenStudy (anonymous):

this is a left sided triangle.

OpenStudy (dumbcow):

i dont know what that means? never heard of left triangle right,equilateral,isosceles, scalene ?

OpenStudy (anonymous):

the triangle in my problem is left sided.

OpenStudy (anonymous):

you need to find out what type of triangle it is. left sided is not one of them.

OpenStudy (dumbcow):

does it have a 90 degree angle

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

it is a "right" triangle hence the answer to your question is above.

OpenStudy (dumbcow):

agreed :)

OpenStudy (anonymous):

yes.